23. The Kp of magnesium hydroxide is 1.2x10-11. Find the concentration of each ion at saturation point.
Answer #1

Mg(OH)2(s) <----------------> Mg+2(aq) + 2OH-(aq)  

Ksp = [Mg+2][OH-]2

Ksp = 1.20 x 10-11

[Mg+2] = S mol /L

[OH-] = 2S mol /L

1.20 x 10-11 = [S] [2S]2

1.20 x1 0-11 = 4S3

S3 = 3.0 x 10-12

S = 6.70 x 10-5 mol/L

[Mg+2] = S = 6.70 x 10-5 mol/L

[OH-] = 2S = 2 x 6.70 x 10-5 = 1.34 x 10-4 mol/L